DAA's stock is selling for $15 per share. The firm's income, assets, and stock price have been growing at an annual 15 percent rate and are expected to continue to grow at this rate for 3 more years. No dividends have been declared as yet, but the firm intends to declare a dividend of D3 = $2.00 at the end of the last year of its fast growth. After that, dividends are expected to grow at a constant growth rate of 6 percent. The firm's required rate of return is 18 percent. Is the stock overpriced or underpriced? By how much?

SHOW ALL CALCULATIONS AND FORMULA IN EXCEL ONLY
Answer #1

We need to use dividend growth model for the same:

Dividend in year 1 and 2 shall be zero as no dividends are paid

YEar 4 dividend =$2.00*1.06=$2.12

Year Div Terminal value Total Discount factor Present Value
1 $          -   $          -   0.847457627 $                      -  
2 $          -   $          -   0.71818443 $                      -  
3 $ 2.0000 $                17.67 $   19.67 0.608630873 $               11.97
4 $      2.12 Total $          11.9697

Discount factors (DF)are calculated as:1/(1+r)^n

Where r is required return and n is the year

For the perpetual growth period, terminal value (TV) of dividends = Dividend next period/(r-g) where r is the required return and g is the growth rate.

Terminal value of dividends= 2.12/(0.18-0.06)=$17.67

As we can see that the stock valued on dividend basis is worth less than $15, hence it is overpriced
Answer #2
Overpriced and $3.03
source: Other web
answered by: Iqra